OpenStudy (anonymous):

Solving System of Equations by Sbstitution. y= -2 4x-3y = 18

sam (.sam.):

y= -2...............(1) 4x-3y = 18..........(2) Substitute (1) to (2), 4x-3y=18 4x-3(-2)=18 4x+6=18 4x=12 x=3